What is the average power dissipation in an RLC series circuit with R=10 ohm, L=0.1 H, C=10 MF, when driven at resonance by 100 V-RMS source


1
Expert's answer
2021-02-09T07:06:58-0500

"P=I_{rms}\times U_{rms}" ; "Z=\sqrt{R^2+(\frac{L}{\sqrt{LC}}-\frac{\sqrt{LC}}{C})^2}"

"I_{rms}=\frac{U_{rms}}{Z}=" "\frac{U_{rms}}{\sqrt{R^2+(\frac{L}{\sqrt{LC}}-\frac{\sqrt{LC}}{C})}}" "=\frac{100}{\sqrt{10^2+(\frac{0.1}{\sqrt{0.1\times10^{-2}}}-\frac{\sqrt{0.1\times10^{-2}}}{10^{-2}})^2}}=10" A

"P=10\times100=1000" W.